The sum of n terms of an ap is n^2 then find the common difference

Grade:11

1 Answers

Arun
6 years ago
Dear Ronit
 
S1 = T1 = 1² = 1
 
S2 = T1 + T2 = 2² = 4
 
S3 = T1 + T2 + T3 = 3² = 9
 
T2 = S2 - S1 = 3
 
Hence T2 - T1 = 3-1 = 2
 
T3 = S3 - S2 = 9-4 = 5
 
T3 - T2 = 2
 
Hence this will form an A.P. with common difference d = 2
 
 
Regards
Arun (askIITians forum expert)